# 03F40 (primary), 03D20

### v2

@@
Abstract
The original Goodstein process is based on writing numbers in hereditary $b$-exponential normal form: that is, each number $n$ is written in some base $b\geq 2$ as $n=b^{e}a+r$, with $e$ and $r$ iteratively being written in hereditary $b$-exponential normal form. We define a new process which generalises the original by writing expressions in terms of a hierarchy of bases $B$, instead of a single base $b$. In particular, the ‘digit’ $a$ may itself be written with respect to a smaller base $b^{\prime}$. We show that this new process always terminates, but termination is independent of Kripke-Platek set theory, or other theories of Bachmann-Howard strength.
1 Introduction
@@
Several recent results have extended Goodstein’s theorem in various ways [1, 9, 10], including variants independent of Kripke-Platek set theory ($\sf KP$) or other theories of Bachmann-Howard strength [2, 11]. The latter are themselves based on functions not provably total in $\sf PA$, with the general philosophy being that they are fair game, as long as they are provably total in $\sf KP$. But this begs the question: are such functions really necessary for Goodstein processes of high proof-theoretic strength? May it be possible to squeeze some more power out of Goodstein’s original construction? A partial answer is ‘no’, because the authors [10] have shown that no way of writing numbers in hereditary exponential normal form will lead to termination times larger than those given by Goodstein’s original progression. However, there is a way to circumvent this restriction.
Our first observation is that there is no need to always increase the base by one; the next base may be chosen to be much larger than the previous one. However, simply replacing the base $b$ by some uniform $f(b)$ will not yield a principle of higher proof-theoretic strength, since the well-foundedness of $\varepsilon_{0}$ may still be applied. Our second observation is that the new base for each $n\in\mathbb{N}$ could depend on $n$. We may consider, say, replacing the base of $n$ by $n$ itself, like the ouroborous eating its own tail. Some reflection shows that this approach will not lead to a uniform base being used at the next step, since following this rule would lead us to replacing the base of $n+1$ by $n+1$, and thus every natural number would be regarded as a possible base at each stage. This naïve idea would not quite work out, but we can follow an intermediate approach and create new bases only at ‘critical’ numbers. This will lead to a hierarchy of bases at each stage, rather than a single ‘global’ base. One can even imagine that for a number written as $b^{e}a+r$ in base $b$, the numbers $b,e,a,r$ may all be written in distinct bases, with their digits written in terms of even smaller bases, each performing its own Goodstein process, like a self-similar fractal made up of smaller copies of itself. As we will see, this can be implemented in a manner that yields a terminating sequence; in this way, we obtain a Goodstein principle very close to the classic version, but of much higher proof-theoretic strength.
@@
2 Fractal Goodstein processes
Let us recall some notions from the classic Goodstein process. Letting $b\in\mathbb{N}\setminus\{0,1\}$ and $0<n\in\mathbb{N}$, we define the $b$-decomposition of $n$ to be $b^{e}a+r$, where $a,e,r$ are uniquely chosen so that $0<a<b$ and $r<b^{e}$, in which case we write $n=_{b}b^{e}a+r$. Technically, we think of the $b$-decomposition of $n$ as the quadruple $(b,e,a,r)\in\mathbb{N}^{4}$. One could iteratively write $e$ and $r$ using their own $b$-decompositions, leading to Goodstein’s hereditary exponential normal forms. Note that $_{b}$ is not symmetric, i.e. $\tilde{b}^{\tilde{e}}\tilde{a}+\tilde{r}=_{B}b^{e}a+r$ does not imply $b^{e}a+r=_{b}\tilde{b}^{\tilde{e}}\tilde{a}+\tilde{r}$; only the right-hand side needs to be a $b$-decomposition. The number $a$ is regarded as a ‘digit,’ since it is less than the base. Formally, for $b\geq 2$, we say that $c$ is a $b$-digit of $n$ if $n=_{b}b^{e}a+r$ and either $c=a$ or $c$ is a $b$-digit of $e$ or $r$. Zero is defined to have no $b$-digits, and if $n<b$, $n=_{b}b^{0}n+0$, so $n$ is the only $b$-digit of $n$. The set of $b$-digits of $n$ is denoted ${\rm Digit}_{b}(n)$.
@@ 2 Fractal Goodstein processes
It is not hard to check that ${\rm base}_{B}(n)={\rm Base}_{B}(n)$, except when $n\in B\setminus\{\min B\}$. For the most part, ${\rm Base}_{B}(n)$ can be regarded as the ‘official’ base of $n$, but ${\rm base}_{B}(n)$ will be useful in Section 4.
Example 2.4
@@
Definition 2.5
@@
Otherwise, $n\geq\min B$. Let $b={\rm Base}_{B}(n)$ and assume inductively that ${\uparrow}m$ is defined for all $m<n$. We first define the deep base-change operator $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle^{C}_{B}$ on $m\in\mathbb{N}$: if $m<b$, then $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m={\uparrow}_{B}^{C}m$. If $m\geq b$, write $m=_{b}b^{e}a+r$, and set
@@
This has the effect of applying the upgrade operator to all the $b$-digits of $m$, while changing the base to $c$.
@@
In words, given $n\in\mathbb{N}$ with $b={\rm Base}_{B}(n)$, in order to perform the upgrade, we first upgrade all the $b$-coefficients of $n$. We then search for the least base $c$ for which changing the base $b$ to $c$ will yield a ‘reasonable’ result: namely, we need $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$ to be at least as large as ${\uparrow}(n-1)$ to make $\uparrow$} monotone, but we also want the result to have $c$ as its base.
When $n$ is a base in $B$, note that $b=n$ and $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n=c$. So, in this case, we basically jump to the next available base in $C$; for this to be a finite value, $C$ should have ‘enough available bases.’ In order for the upgrade to be well-behaved, we also do not want this value to be ‘much larger’ than ${\uparrow}_{B}^{C}(n-1)$. The precise conditions required are given by Definition 2.9 below.
Convention 2.6
If in a result or proof we state that transformations are from $B$ to $C$, this indicates that $\uparrow$} is to be understood as ${\uparrow}^{C}_{B}$ and $\langle\begin{smallmatrix}\cdot\\$ $\cdot\end{smallmatrix}\rangle$ as $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle^{C}_{B}$. Similarly, we may state e.g. that $\uparrow$} is from $B$ to $C$ to indicate that ${\uparrow}={\uparrow}^{C}_{B}$. More often than not, these indices can be suppressed, but there will be occasions when we do consider multiple base hierarchies in the same context.
Convention 2.7
@@
Example 2.8
Consider transformations between $B=\{2,6\}$ and $C=\{3,30,180\}$. We have that ${\uparrow}0=0$ and ${\uparrow}1=1$. Then, ${\rm Base}_{B}(2)=2$ and $\langle\begin{smallmatrix}3\\ 2\end{smallmatrix}\rangle 2=3<30=S_{C}(3)$, so ${\uparrow}2=3$. More generally, when $B$ and $C$ are base hierarchies with $\min C\geq\min B$, ${\uparrow}^{C}_{B}\min B=\min C$.
Now, let $n=6^{6}+6^{2^{2}+1}$ and let us compute ${\uparrow}_{B}^{C}n$. We first perform the upgrade on $2^{2}+1$. Writing $\uparrow$} instead of ${\uparrow}_{B}^{C}$ for simplicity, in this case ${\uparrow}(2^{2}+1)=3^{3}+1<30$, exactly as we would expect from the original Goodstein process.
@@
Example 2.10
Let $B$ be any base hierarchy. We define $B_{+}=\bigcup_{n<\infty}C_{n}$, where $C_{n}$ is defined by induction on $n$ as follows:
  1. 1.
    $C_{n}=\{\min B+1\}$ if $n\leq\min B$,
  2. 2.
    $C_{n+1}=C_{n}$ if $n+1\notin B$, and
  3. 3.
    if $\min B<n+1\in B$, then $C_{n+1}=C_{n}\cup\{k\}$, where $k$ is the least multiple of $\max C_{n}$ strictly above ${\uparrow}_{B}^{C_{n}}n$.
Then, $B_{+}$ is always a good successor for $B$, a fact that will follow from the more general Proposition 5.7 in view of Example 5.3.
As an example, let $B=\{2,6\}$, and let us compute $C:=B_{+}$. First, we see that $C_{n}=\{\min B+1\}=\{3\}$ for $n\leq 3$. From here, we see that ${\uparrow}_{B}^{C_{5}}5=3^{3}+1=28$ and ${\rm Base}_{C_{5}}(10)=3$, so $C_{6}$ is obtained by adding the next multiple of $3$, namely, $30$. We thus arrive at $C=\{3,30\}$, and it is not hard to check that $C$ is indeed a good successor for $B$.
@@
Example 2.13
Let $B_{0}$ be any base hierarchy and define inductively $B_{i+1}=(B_{i})_{+}$, as in Example 2.10. Then, $(B_{i})_{i\in\mathbb{N}}$ is a dynamical hierarchy. If $B_{0}=\{2\}$, we obtain the dynamical hierarchy of Example 2.12 as a special case. Note moreover that if $B_{0}$ is finite, then $B_{i}$ is finite for all $i$.
@@
Definition 2.14
Given a dynamical hierarchy $\mathcal{B}$ and $n\in\mathbb{N}$, we recursively define ${\mathbb{G}}^{\mathcal{B}}_{i}(n)$ by letting ${\mathbb{G}}^{\mathcal{B}}_{0}(n)=n$ and, if ${\mathbb{G}}^{\mathcal{B}}_{i}(n)$ is defined and positive, then ${\mathbb{G}}^{\mathcal{B}}_{i+1}(n)={\uparrow}_{i}{\mathbb{G}}^{\mathcal{B}}_{i}(n)-1$; otherwise ${\mathbb{G}}^{\mathcal{B}}_{i+1}(n)$ is undefined.
Example 2.15
@@
With this, we may state our main results, proven throughout the rest of the paper.
Theorem 2.16
@@
Theorem 2.17
There is a dynamical hierarchy $\mathcal{O}$ such that $\sf KP$ does not prove that for every $n\in\mathbb{N}$, there exists $i\in\mathbb{N}$ with ${\mathbb{G}}^{\mathcal{O}}_{i}(n)=0$.
The dynamical hierarchy $\mathcal{O}$ is computable and will be given by Definition 5.1. Thus Theorem 2.17 should be understood as follows. Say that an arithmetical formula $\delta$ is bounded if all quantifiers occurring in $\delta$ are of the form $\forall u<t(\vec{v})$ or $\exists u<t(\vec{v})$, where $t$ is a polynomial with coefficients in $\mathbb{N}$, and $\varphi$ is a $\Sigma^{0}_{1}$-formula if it is of the form $\exists z\delta$, where $\delta$ is bounded. Arithmetical formulas, and $\Sigma^{0}_{1}$-formulas in particular, can be encoded in the language of $\sf KP$ in a standard way [3]. Since ${\mathbb{G}}^{\mathcal{O}}_{y}(x)$ is computable [18], there is a $\Sigma^{0}_{1}$-formula $\varphi(x,y)$ such that $\varphi(x,y)$ is true on $\mathbb{N}$ if and only if ${\mathbb{G}}^{\mathcal{O}}_{y}(x)=0$. Then, Theorem 2.17 states that $\sf KP$ does not prove $\forall x\in\mathbb{N}\ \exists y\in\mathbb{N}\ \varphi(x,y)$.
3 Properties of the Upgrade Operator
Our first main goal is to prove that the fractal Goodstein process is always terminating. This is based on various regularity properties of the upgrade operator on good successors, which we establish in this section. We begin with what is perhaps the most crucial property, which is that the upgrade is strictly increasing.
Lemma 3.1
Let all transformations be from a base hierarchy $B$ to a good successor $C$, $m<n\in\mathbb{N}$, and $b\geq 2$.
  1. 1.
    ${\uparrow}m<{\uparrow}n$.
  2. 2.
    If $1<c\leq d$, then $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m\leq\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle m<\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$.
Proof 3.2.
Note that we may assume without loss of generality that m=n1m=n-1, since otherwise we can see by induction on $n$ that $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle m\leq\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle{(n-1)}<\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$ and ${\uparrow}m\leq{\uparrow}(n-1)<{\uparrow}n$. The first claim is then immediate, as ${\uparrow}n$ is explicitly selected larger than ${\uparrow}(n-1)$ whenever ${\uparrow}n<\infty$.
The second claim is proven by induction on $n$. If $n<b$, we can apply the first claim to obtain ${\uparrow}m<{\uparrow}n$, and thus

### v4

# The Fractal Goodstein Principle

@@
Abstract.
The original Goodstein process is based on writing numbers in hereditary $b$-exponential normal form: each number $n$ is written in some base $b\geq 2$ as $n=b^{e}a+r$, with $e$ and $r$ iteratively being written in hereditary $b$-exponential normal form. We define a new process which generalises the original by writing expressions in terms of a hierarchy of bases $B$, instead of a single base $b$. In particular, the ‘digit’ $a$ may itself be written with respect to a smaller base $b^{\prime}$. We show that this new process always terminates, but termination is independent of Kripke-Platek set theory, or other theories of Bachmann-Howard strength.
Key words and phrases:
Goodstein’s theorem, Gödel incompleteness, Howard-Bachmann ordinal, fast-growing functions
2020 Mathematics Subject Classification:
Primary 03F40, 03D20, 03D60
The first author was supported by the Spanish Ministry of Science and Innovation grant PID2023-149556NB-I00.
1. Introduction
@@
Several recent results have extended Goodstein’s theorem in various ways [1, 9, 10], including variants independent of Kripke-Platek set theory ($\sf KP$) or other theories of Bachmann-Howard strength [2, 11]. The latter are themselves based on functions not provably total in $\sf PA$, with the general philosophy being that they are fair game, as long as they are provably total in $\sf KP$. But this begs the question: are such functions really necessary for Goodstein processes of high proof-theoretic strength? May it be possible to squeeze some more power out of Goodstein’s original construction? A partial answer is ‘no’, because the authors [10] have shown that no way of writing numbers using elementary functions will lead to termination times larger than those given by Goodstein’s original progression. However, there is a way to circumvent this restriction.
Our first observation is that there is no need to always increase the base by one; the next base may be chosen to be much larger than the previous one. However, simply replacing the base $b$ by some uniform $f(b)$ will not yield a principle of higher proof-theoretic strength, since the well-foundedness of $\varepsilon_{0}$ may still be applied. This leads to our second observation: the new base for each $n\in\mathbb{N}$ could depend on $n$. We may consider, say, replacing the base of $n$ by $n$ itself. Such an approach will not lead to a uniform base being used at the next step, since following this rule would lead us to replacing the base of $n+1$ by $n+1$, and thus every natural number would be regarded as a possible base at each stage. This naïve idea would not quite work out, but we can follow an intermediate approach and create new bases only at ‘critical’ numbers. This will lead to a hierarchy of bases at each stage, rather than a single ‘global’ base. One can even imagine that for a number written as $b^{e}a+r$ in base $b$, the numbers $b,e,a,r$ may all be written in distinct bases, with their digits written in terms of even smaller bases, each performing its own Goodstein process, like a self-similar fractal made up of smaller copies of itself. As we will see, this can be implemented in a manner that yields a terminating sequence; in this way, we obtain a Goodstein principle very close to the classic version, but of much higher proof-theoretic strength.
@@
2. Fractal Goodstein Processes
Let us recall some notions from the classic Goodstein process. Letting $b\in\mathbb{N}\setminus\{0,1\}$ and $0<n\in\mathbb{N}$, we define the $b$-decomposition of $n$ to be $b^{e}a+r$, where $a,e,r$ are uniquely chosen so that $0<a<b$ and $r<b^{e}$, in which case we write $n=_{b}b^{e}a+r$. Technically, we think of the $b$-decomposition of $n$ as the quadruple $(b,e,a,r)\in\mathbb{N}^{4}$, but one could iteratively write $e$ and $r$ using their own $b$-decompositions, leading to Goodstein’s hereditary exponential normal forms. Note that $_{b}$ is not symmetric, i.e. $\tilde{b}^{\tilde{e}}\tilde{a}+\tilde{r}=_{B}b^{e}a+r$ does not imply $b^{e}a+r=_{b}\tilde{b}^{\tilde{e}}\tilde{a}+\tilde{r}$; only the right-hand side needs to be a $b$-decomposition. The number $a$ is regarded as a ‘digit,’ since it is less than the base. Formally, for $b>0$, we say that $c$ is a $b$-digit of $n$ if $n=_{b}b^{e}a+r$, and either $c=a$, or $c$ is a $b$-digit of $e$ or $r$. Zero is defined to have no $b$-digits, and if $0<n<b$, $n=_{b}b^{0}n+0$, so $n$ is the only $b$-digit of $n$. The set of $b$-digits of $n$ is denoted ${\rm Digit}_{b}(n)$.
@@ 2. Fractal Goodstein Processes
It is not hard to check that ${\rm base}_{B}(n)={\rm Base}_{B}(n)$, except when $n\in B\setminus\{\min B\}$. For the most part, ${\rm Base}_{B}(n)$ can be regarded as the ‘official’ base of $n$, but ${\rm base}_{B}(n)$ will be useful in Section 4 for proving that our Goodstein process is terminating.
Example 2.4.
@@
Our Goodstein principle is based on the upgrade operator ${\uparrow}={\uparrow}_{B}^{C}$, which takes the role of $\left(\begin{smallmatrix}c\\ b\end{smallmatrix}\right)$ in Goodstein’s original theorem. The purpose of this operator is to transform a number $n$ represented with respect to a base hierarchy $B$ to a new number ${\uparrow}n$ represented with respect to another base hierarchy $C$. Intuitively, given $n\in\mathbb{N}$ with $b={\rm Base}_{B}(n)$, in order to perform the upgrade, we first recursively upgrade all the $b$-digits of $n$. We then search for the least base $c$ for which changing the base $b$ to $c$ will yield a ‘reasonable’ result: namely, we need ${\uparrow}n$ to be larger than ${\uparrow}(n-1)$ to make $\uparrow$} monotone, but we also want the result to have $c$ as its base, so the end result should be smaller than $S_{C}(c)$. This process of upgrading the digits first and then changing the base $b$ of $n$ to some $c\in C$ is the deep base change, and the result is denoted by $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$. Let us make these operations precise.
Definition 2.5.
@@
Otherwise, $n\geq\min B$. Let $b={\rm Base}_{B}(n)$ and assume inductively that ${\uparrow}m$ is defined for all $m<n$. We first define the deep base-change operator $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle^{C}_{B}$ on $m\in\mathbb{N}$: if $m<b$, then $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m={\uparrow}m$. If $m\geq b$, write $m=_{b}b^{e}a+r$, and set
@@
This has the effect of applying the upgrade operator to all the $b$-digits of $m$, then changing the base to $c$.
@@
When $n$ is a base in $B$, note that $b=n$ and $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n=c$. So, in this case, we basically jump to the next available base in $C$; for this to be a finite value, $C$ should have ‘enough available bases.’ In order for the upgrade to be well-behaved, we also do not want this value to be ‘much larger’ than ${\uparrow}(n-1)$. The precise conditions required from $C$ are given by Definition 2.9 below.
Convention 2.6.
If in a result or proof we state that transformations are from $B$ to $C$, this indicates that $\uparrow$} is to be understood as ${\uparrow}^{C}_{B}$ and $\langle\begin{smallmatrix}\cdot\\$ $\cdot\end{smallmatrix}\rangle$ as $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle^{C}_{B}$. Similarly, we may state e.g. that $\uparrow$} is from $B$ to $C$ to indicate that ${\uparrow}={\uparrow}^{C}_{B}$. More often than not, these indices can be suppressed, but there will be occasions when we do consider more than two base hierarchies in the same context.
Convention 2.7.
@@
Example 2.8.
Consider transformations between $B=\{2,6\}$ and $C=\{3,30,180\}$ and let us compute some values of the upgrade operator. We have that ${\uparrow}0=0$ and ${\uparrow}1=1$. Then, ${\rm Base}_{B}(2)=2$ and $\langle\begin{smallmatrix}3\\ 2\end{smallmatrix}\rangle 2=3<30=S_{C}(3)$, so ${\uparrow}2=3$. More generally, when $B$ and $C$ are base hierarchies with $\min C\geq\min B$, ${\uparrow}^{C}_{B}\min B=\min C$.
Now, let $n=6^{6}+6^{2^{2}+1}$ and let us compute ${\uparrow}_{n}$. We first perform the upgrade on $2^{2}+1$. In this case, ${\uparrow}(2^{2}+1)=3^{3}+1<30$, exactly as we would expect from the original Goodstein process.
@@
In other words, a good successor is one for which the upgrade operator is well-defined and bases always jump to the least available multiple of the previous base. Our main results will involve Goodstein processes defined on sequences of good successors, so it is a good idea to check that these actually exist. Below, we will define a rather general construction to build good successors of arbitrary bases. The intuition is that good successors are typically constructed iteratively, in tandem with $\uparrow$}. That is, if $B$ is a base hierarchy, we build a good successor as an increasing union of finite sets $C=\bigcup_{n\in\mathbb{N}}C_{n}$, in such a way that ${\uparrow}_{B}^{C}x={\uparrow}_{B}^{C_{n}}x$ whenever $x\leq n$. If $n+1\in B$, we add a new base to $C_{n+1}$ to ensure that the good successor condition holds up to $n+1$. We may also add new bases in cases where $n+1\notin B$ if we desire, but in order for the upgrade to behave nicely, it is convenient to only do so when $n+1$ is a multiple of its own base. Such a number is $B$-critical, and the set of $B$-critical numbers is denoted $\partial B$. The picture is then that we choose some $X$ with $B\subseteq X\subseteq\partial B$, and add a new base to $C_{n}$ only when $n\in X$. Let us make this precise.
Definition 2.10.
Let $B$ be a base hierarchy. Say that $n>0$ is $B$-critical if ${\rm Base}_{B}(n)\mid n$. Let $\partial B$ denote the set of $B$-critical numbers.
Given $X$ with $B\subseteq X\subseteq\partial B$, we define a new base hierarchy $B^{X}=\bigcup_{n<\infty}B^{X}_{n}$. Let $n\in\mathbb{N}$ and assume inductively that $B^{X}_{m}$ has been defined and is finite for all $m<n$. Then, we define $B^{X}_{n}$ according to the following cases.
  1. Case 1
    ($n\leq\min B$). $B^{X}_{n}=\{\min B+1\}$.
  2. Case 2
    ($\min B<n\notin X$). $B^{X}_{n}=B^{X}_{n-1}$.
  3. Case 3
    ($\min B<n\in X$). Let $c=\max B^{X}_{n-1}$ and transformations be from $B$ to $B^{X}_{n-1}$. Let $a$ be least so that ${\uparrow}(n-1)<ca$. Then, set [^1] As a failsafe, we may set $B^{X}_{n}=B^{X}_{n-1}$ if ${\uparrow}(n-1)=\infty$; however, as we will see, this will never be the case. $B^{X}_{n}=B^{X}_{n-1}\cup\big\{ca\}$.
If $X=B$, we define $B^{\prime}_{n}=B^{X}_{n}$ and $B^{\prime}=B^{X}$. This is the minimalistic successor of $B$. Likewise, if $X=\partial B$, we define $B^{+}_{n}=B^{X}_{n}$ and $B^{+}=B^{X}$ and call this the greedy successor of $B$.
The new base hierarchy $B^{X}$ is always a good successor for $B$. We defer the proof until Proposition 5.2, as this requires some machinery. Note that the map $X\mapsto B^{X}$ is not monotone, so $B^{\prime}$ is not minimal under set inclusion. It is, however, the ‘leftmost’ good successor of the form $B^{X}$, if we identify bases with binary reals $(r_{i})_{i\in\mathbb{N}}$, with $r_{i}=1$ if $i\in B$, otherwise, $r_{i}=0$. Similarly, $B^{+}$ is the rightmost successor that is of the form $B^{X}$.
Remark 2.11.
Good successors of the form $B^{X}$ are by no means exhaustive: for example, we can let $\min C>\min B+1$ but otherwise follow Definition 2.10 to obtain a good successor to the left of $B^{\prime}$. Likewise, we can modify the case for $n\in X$ to add any finite number of bases $ca,2ca,4ca,\ldots$ and obtain a good successor which is to the right of $B^{+}$. However, good successors of the form $B^{X}$ (or even just $B^{+}$) will suffice for establishing our main results, so we will not explore other constructions further in this text.
Example 2.12.
Let $B=\{2,6\}$, and let us compute $B^{\prime}$. First, we see that $B^{\prime}_{n}=\{\min B+1\}=\{3\}$ for $n\leq 5$. From here, we see that ${\uparrow}_{B}^{B^{\prime}_{5}}5=3^{3}+1=28$ and ${\rm Base}_{B^{\prime}_{5}}(28)=3$, so $B^{\prime}_{6}$ is obtained by adding the next multiple of $3$, namely, $30$. We thus arrive at $B^{\prime}=\{3,30\}$, and it is not hard to check that $B^{\prime}$ is indeed a good successor for $B$.
Example 2.13.
Let $B=\{2\}$ and let us compute $B^{+}_{6}$. Similarly to Example 2.12, $B^{+}_{3}=\{3\}$, but now, to find $B^{+}_{4}$, we first compute ${\uparrow}_{B}^{B^{+}_{3}}3={\uparrow}_{B}^{B^{+}_{3}}(2+1)=4$. Since ${\rm Base}_{B^{+}_{3}}(4)=3$, we must add the next multiple of $3$ after $4$, that is, $6$. We thus arrive at $B^{+}_{4}=B^{+}_{3}\cup\{6\}=\{3,6\}$. Since $2\nmid 5\notin B$, $B^{+}_{5}=B^{+}_{4}$. On the other hand, we do add a new element to $B^{+}_{6}$. Compute ${\uparrow}_{B}^{B^{+}_{5}}5={\uparrow}_{B}^{B^{+}_{5}}(2^{2}+1)=6^{6}+1$. The next multiple of $6$ is $6^{6}+6$, so $B^{+}_{6}=B^{+}_{5}\cup\{6^{6}+6\}=\{3,6,6^{6}+6\}$.
In general, we will add one new base for every even number, and thus $B^{+}$ is infinite, even though $B$ has only one element.
@@ 3. Case 3
Example 2.16.
Let $B_{0}$ be any base hierarchy and define inductively $B_{i+1}=(B^{\prime})_{i}$, as in Example 2.12. Then, $(B_{i})_{i\in\mathbb{N}}$ is a dynamical hierarchy. If $B_{0}=\{2\}$, we obtain the dynamical hierarchy of Example 2.15 as a special case. Note moreover that if $B_{0}$ is finite, then $B_{i}$ is finite for all $i$.
Our independence results will be based on a specific dynamical hierarchy built using greedy successors.
Definition 2.17.
Let $\mathcal{D}=(\mathcal{D}_{i})_{i\in\mathbb{N}}$ be given by $\mathcal{D}_{0}=\{2\}$ and $\mathcal{D}_{i+1}=\mathcal{D}^{+}_{i}$; this is the greedy dynamical hierarchy.
@@
Definition 2.18.
Given a dynamical hierarchy $\mathcal{B}$ and $n\in\mathbb{N}$, we recursively define ${\mathbb{G}}^{\mathcal{B}}_{i}(n)$ by letting ${\mathbb{G}}^{\mathcal{B}}_{0}(n)=n$ and, if ${\mathbb{G}}^{\mathcal{B}}_{i}(n)$ is defined and positive, then ${\mathbb{G}}^{\mathcal{B}}_{i+1}(n)={\uparrow}_{i}{\mathbb{G}}^{\mathcal{B}}_{i}(n)-1$; otherwise, ${\mathbb{G}}^{\mathcal{B}}_{i+1}(n)$ is undefined.
Example 2.19.
@@
With this, we may state our main results, proven throughout the rest of the text.
Theorem 1.
@@
Theorem 2.
Let $\mathcal{D}$ be the greedy dynamical hierarchy of Definition 2.17. Then, $\sf KP$ does not prove that for every $n\in\mathbb{N}$, there exists $i\in\mathbb{N}$ with ${\mathbb{G}}^{\mathcal{D}}_{i}(n)=0$.
Theorem 2 should be understood as follows. Say that an arithmetical formula $\delta$ is bounded if all quantifiers occurring in $\delta$ are of the form $\forall u<t(\vec{v})$ or $\exists u<t(\vec{v})$, where $t$ is a polynomial with coefficients in $\mathbb{N}$, and $\varphi$ is a $\Sigma^{0}_{1}$-formula if it is of the form $\exists z\in\mathbb{N}\ \delta(z)$, where $\delta(z)$ is bounded. Arithmetical formulas, and $\Sigma^{0}_{1}$-formulas in particular, can be encoded in the language of $\sf KP$ in a standard way [3]. Since ${\mathbb{G}}^{\mathcal{D}}_{y}(x)$ is computable [19], there is a $\Sigma^{0}_{1}$-formula $\varphi(x,y)$ such that $\varphi(x,y)$ is true on $\mathbb{N}$ if and only if ${\mathbb{G}}^{\mathcal{D}}_{y}(x)=0$. Then, Theorem 2 states that $\sf KP$ does not prove $\forall x\in\mathbb{N}\ \exists y\in\mathbb{N}\ \varphi(x,y)$.
3. Properties of the Upgrade Operator
The proofs of our main results rely heavily on various regularity properties of the upgrade operator and the deep base change, which we establish in this section. We begin with what is perhaps their most fundamental property: monotonicity.
Lemma 3.1.
Let all transformations be from a base hierarchy $B$ to a good successor $C$ and $m,n,b,c,d\in\mathbb{N}$ with $b,c,d\geq 2$.
  1. (1)
    If $m<n$, then ${\uparrow}m<{\uparrow}n$.
  2. (2)
    If $m<n$ and ${\uparrow}b\leq c$, then $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle m\leq\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$.
  3. (3)
    If $c<d$, then $\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n\leq\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$, and the inequality is strict if $n\geq b$.
  4. (4)
    If $b<n$, then $c<\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n$, and $c=\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle b$.
Proof.
(1) Note that we may assume without loss of generality that m=n1m=n-1, since otherwise we can see by induction on $n$ that ${\uparrow}m\leq{\uparrow}(n-1)<{\uparrow}n$. The claim is then immediate, as ${\uparrow}n$ is explicitly selected larger than ${\uparrow}(n-1)$ whenever ${\uparrow}n<\infty$.
(2) As above, assume that m=n1m=n-1 and proceed by induction on $n$. If $n<b$, we can apply the first claim to obtain ${\uparrow}m<{\uparrow}n$, and thus